OpenStudy (anonymous):

Solve the system of following linear equations by addition: -4x+y=-4 8x+y=20

OpenStudy (rulnick):

from first eqn: y=4x-4 subst in second eqn: 8x+4x-4=20 solve 12x=24 x=2 back substitute to get y: y=4(2)-4 = 4. So x=2, y=4. Done.

OpenStudy (anonymous):

Where did you get the equation to back substitute in?

OpenStudy (rulnick):

y=4x-4 with x=2